What is the PH at the equivalence point in the titration of 100ml of 0.20m ammonia with 0.10M hydrochloric acid?

Given NH3 as 1.75×10^-5


1
Expert's answer
2020-12-01T09:20:14-0500

NH3 + HCl = NH4+ + Cl-

Moles of NH3 = 0.2 × 100 ×10-3 = 0.02 moles

Vol HCl = 0.02moles/ 0.10M = 0.2L

Total volume = 0.2+ 0.1 = 0.3L

Concentration of ammonium ion = 0.02/0.3 = 0.067M

Ka = Kw/Kb = 10-14/ 1.75×10-5 = 5.71 × 10-10

Ka = NH3 . H3O / NH4+

= x*x/(0.067-x)

0.067-x ~ 0.067

Ka= x²/ 0.067 = 5.71 × 10-10

x= (0.067 × 5.71 ×10-10)1/2

x= 6.19 × 10-6 = [H3O]+

PH = -log [6.19 ×10-6]

= 5.21